OpenStudy (anonymous):

Thanks for helping! If I have the coordinates (0,7) and I need to use the formula tan theta=y/x, (x cannot equal 0), how do I find the angle? Isn't it undefined? Thank You!

OpenStudy (amistre64):

there is only one angle ... well two ... that have an undefined tangent, but then again their cotangent IS defined. what are they?

OpenStudy (anonymous):

I'm not sure :(

OpenStudy (amistre64):

think of it this way: tan = sin/cos for what angle is cos(t) = 0?

OpenStudy (anonymous):

For sin(t)=0?

OpenStudy (amistre64):

no, its the bottom that is zero ... so cos(t) = 0

OpenStudy (amistre64):

y/x = sin(a)/cos(a) y cos(a) = x sin(a) since x=0 and y=7 7 cos(a) = 0 cos(a) = 0

OpenStudy (anonymous):

ah, so it is pi/2?

OpenStudy (amistre64):

yes :)

OpenStudy (anonymous):

awesome, that was really helpful!

OpenStudy (anonymous):

Thank You!

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

:D